iosangel:

Larry's Lemons is a street vendor business that sells lemonade and lemon bars. A cup of lemonade sells for $2 and a lemon bar sells for $1.50. When all related business expenses are included, a cup of lemonade costs $0.25 to prepare and a lemon bar costs $0.20 to prepare. Last Monday, one of the vendors selling Larry's Lemons sold at least $500 worth of lemonade and lemon bars and its expenses were no more than $100. At least 150 cups of lemonade were sold. Let x be the number of cups of lemonade sold last Monday and y be the number of lemon bars sold last Monday. Which ordered pairs representing a combination of cups of lemonade and lemon bars could have been sold last Monday and make sense in the context of the situation? Select each correct answer. (160,110) (232.5,200) (155,305.5) (150,200) (180,100)

Bob:

2x + 1.5y = 500 0.25x + 0.2y = 100 These should be your equations, do you know what to do next

Bob:

@iosangel

iosangel:

solve?

Bob:

do you know how to solve

iosangel:

substitiion?

Bob:

no it's simple, just plug in the x values in the choices for x, and do the same for Y

iosangel:

i meant like substitiion method

Bob:

for the first equation, try the first choice first, and it has to match the 2nd equation. 2(160) + 1.5(110) = 500

Bob:

do you understand

iosangel:

x=−2000 and y=3000

Bob:

oof I might be wrong you know what I'll tag somebody else @Shadow

iosangel:

ok thank you